OpenStudy (anonymous):

Simplify (answer using positive exponents only): (a^-3 x b / c^-2)^-4 In other words, a-3 x b over c^-2, all powered to -4. [I seriously cannot figure out how to write equations here. Gah.]

OpenStudy (anonymous):

a^3 / c^2b all raised to 4 power

OpenStudy (mertsj):

\[(\frac{a ^{-3}b}{c ^{-2}})^{-4}=\frac{a ^{12}b ^{-4}}{c ^{8}}=\frac{a ^{12}}{b ^{4}c ^{8}}\]

OpenStudy (mertsj):

Remember...when you raise to a power, you multiply the exponents.

OpenStudy (anonymous):

Thank you! My cousin and I were confused with what to do first: the -4, or flipping the variables.

OpenStudy (mertsj):

Well it really doesn't matter that much but in the end you want the exponents to be positive so do whichever will facilitate the fewest steps and therefore less work.